# Evaluating Privacy Perceptions, Experience, and Behavior of Software Development Teams

### v1

@@ 9 Conclusion
In this paper, we examined privacy perceptions, practices, and awareness of SDLC team members during software development. Our findings suggest a need for standardized privacy practices, educational awareness and implementation of PbD, and a privacy expert to promote privacy awareness and compliance. We identified gaps in privacy practices among software teams. Finally, we provide research and educational directions to reduce the challenges in implementing these practices.
@@
Appendix A Participants’ Demographic Information
@@
Appendix B Details of Solove’s Taxonomy
@@
Appendix C Confidence in Security & Privacy Measures
@@
Appendix D Presence of a CPO or a Similar Role
@@
The list of the hypotheses for the correlation between the presence of a CPO or a similar role and the creation of a PIA, familiarity with PETs, number of privacy breaches, and the size of a company are as follows:
@@
Appendix E The Creation of a PIA
The list of the hypotheses for the correlation between the creation of a PIA and the size of a company and confidence in privacy and security measures are as follows:
@@ Appendix E The Creation of a PIA
The distribution of responses to the creation of a PIA in their company is shown below.
@@
Appendix F Privacy by Design Approaches
@@
Appendix G The Most Commonly Used PETs
Data Protection and Encryption: A widely adopted technique, encryption converts data into a code to prevent unauthorized access. It is used for both data at rest (stored data) and data in transit. Examples of encryption techniques mentioned include AES-256 [51] and SSL/TLS [50].
Access Control: This involves implementing policies that restrict access to sensitive data to authorized individuals only. Role-based access control, where users are granted access to specific features based on their roles in the organization, is a common practice [52].
Anonymity and Pseudonymity: These techniques involve removing personally identifiable information (PII) to prevent individual users from being identified. Pseudonymity involves replacing identifiers with pseudonyms [60].
Differential Privacy Approaches: This technique involves adding noise to the data to protect individual user information while still providing useful insights. It is particularly useful in data analysis and machine learning applications [16].
Secure Communication: This involves encrypting all communications within the software using standard protocols like HTTPS and SSL/TLS [50].
Privacy-Enhanced Web Browsing: This method uses technologies like tracker blocking, web request encryption, and privacy-centric search engines to enhance online privacy. It minimizes the personal data third parties can gather during web browsing.
Privacy-Focused Hardware: This refers to hardware built with privacy at its core, featuring physical switches to disable microphones and cameras, and often running on open-source software. It aims to give users control and protect their data from unauthorized access.
Privacy Governance and Compliance Technologies: These systems help organizations comply with data protection laws like GDPR and HIPAA by assessing, managing, and reporting data privacy practices.
Privacy-Focused Operating Systems and Tools: They are built to prioritize user privacy and reduce data exposure. They usually feature pre-installed privacy tools and default encryption. They might include messaging apps with end-to-end encryption or services that avoid tracking user data.
@@ Appendix I Factors Influencing Usage of Forums
As shown in Table 19 when comparing forum usage with the size of the company, a statistically significant difference was found between the groups ($H-Value=4.03,p-value=0.04$). However, no significant difference was noted when comparing forum usage with the presence of a Chief Privacy Officer (CPO) ($H-value=9.83,p-value=0.08$). No significant differences were detected when comparing forum usage with the usage of Privacy Enhancing Technologies (PETs) ($H-value=3.92,p-value=0.56$). These findings suggest that the size of the company and the presence of a CPO are more likely to influence the frequency with which developers consult forums for privacy-related inquiries, while the use of PETs appears to have less impact on this behavior.
Table 19: P-Value and H Value for Hypothesis H5a to H5c
H5a H5b H5c
P-Value 0.04 0.17 0.08
H Value 4.03 7.80 9.83
@@
Below is the list of hypotheses for location analysis. and

### v2

@@ 9 Conclusion
In this paper, we examined privacy perceptions, practices, and behaviors of SDLC team members during software development. Our findings suggest a need for standardized privacy practices, educational awareness and implementation of PbD, and a privacy expert to promote privacy awareness and compliance. We identified gaps in privacy practices among software teams. Finally, we provide research and educational directions to reduce the challenges in implementing these practices.
@@
Acknowledgments
This research was funded by NSF Award # 2238047.
Appendix A Survey Questions
Survey questions can be found here: http://tinyurl.com/2p9n49e4
Appendix B Participants’ Demographic Information
@@
Appendix C Details of Solove’s Taxonomy
@@
Appendix D Confidence in Security & Privacy Measures
@@
Appendix E Presence of a CPO or a Similar Role
@@
The hypotheses list for the correlation between the presence of a CPO/a similar role and the PIA creation, familiarity with PETs, number of privacy breaches, and the company size are:
@@
Appendix F The Creation of a PIA
The hypotheses list for the correlation between the creation of a PIA and the company size and confidence in privacy and security measures are:
@@ Appendix F The Creation of a PIA
The distribution of responses to the creation of a PIA in their company is shown in Figure 5.
@@
Appendix G Privacy by Design Approaches
@@ Appendix H Detailed Analysis of PETs’ Familiarity
Table 19: P-Value and H Value for Hypothesis H5a to H5c
H5a H5b H5c
P-Value 0.04 0.17 0.08
H Value 4.03 7.80 9.83
Table 20: Categories of PETs
Categories of PETs Definition in Literature [62, 76] Example from Our Survey
Encryption A system of communication where the only people who can read the messages are the people communicating. We use encryption and a number of security features offered by the platform we implement. It is primarily the responsibility of the back-end programmers.
Access Control/ Identity protection Deals with identifying individuals and controlling access to resources in a system. We implement role-based access for the various features of our product as well as internally
Anonymity and Pseudonymity Involves removing personally identifiable information (PII) to prevent individual users from being identified. Pseudonymity involves replacing identifiers with pseudonyms [83]. Data anonymization, our managers would be the primary users for that subject
Differential Privacy Involves adding noise to the data to protect individual user information while still providing useful insights. It is particularly useful in data analysis and machine learning applications. [25] We use encryption and al little bit of differential privacy where it is applicable and it varies from project to project with who is tasked with implementing these features.
Secure Communication/ VPN Involves encrypting all communications within the software using standard protocols like HTTPS and SSL/TLS. All of our internal communication is done over an internal VPN, and all web access is done with https.
Privacy-Enhance Anti Web Tracking Involves blocking attempts of different types of trackers to monitor users’ online activity and personal data. -
@@ Appendix I Factors Influencing Usage of Forums
As shown in Table 19 when comparing forum usage with the size of the company, a statistically significant difference was found between the groups ($H-Value=4.03,p-value=0.04$). However, no significant difference was noted when comparing forum usage with the presence of a Chief Privacy Officer (CPO) ($H-value=9.83,p-value=0.08$) or with the usage of Privacy Enhancing Technologies (PETs) ($H-value=3.92,p-value=0.56$). These findings suggest that only the size of the company is more likely to influence the frequency with which developers consult forums for privacy-related inquiries.
@@
Below is the list of hypotheses for location analysis.
@@
Appendix K Qualitative Analysis Guidelines
Table 20 shows the different categories of PETs and Table 21 describes the privacy taxonomy, both of which were considered as guidelines for our qualitative analysis.
Table 21: Taxonomy of Privacy
Taxonomy of Privacy Solove’s Definition [79] Example from IAPP [50] Example from Our Survey
Surveillance Watching, listening to, or recording of an individual’s activities A website monitoring the cursor movements of a visitor while visiting the website. Privacy is the ability to keep information or activities out of public knowledge
Interrogation Questioning or probing for personal information An interviewer asking an inappropriate question, such as marital status, during a employment interview. As far as I’m the internet, not asking for private information from our customers such as addresses or any sensitive information.
Aggregation Combining of various pieces of personal information A credit bureau combining an individual’s payment history from multiple creditors. Keeping unnecessary information from being exchanged at the minimum amount possible.
Insecurity Carelessness in protecting information from leaks or improper access An e-commerce website allowing others to view an individual’s purchase history by changing the URL (e.g. enterprivacy.com?id=123) Having confidential and private information secured and stored away safely from malicious users.
Identification Linking of information to a particular Individual. A researcher linking medical files to the Governor of a state using only date of birth, zip code and gender. I think it can be defined as a set of personal information of each individual that should not be accessible to other people
Secondary Use Using personal information for a purpose other than the purpose or which is was collected The U.S. Government uses census data collected for the purpose of apportioning Congressional districts to identify and intern those of Japanese descent in WWII. Ensuring the minimum amount of data is available only to those that genuinely need it for business purposes, and that it’s only available for the specified amount of time that the data is needed.
Exclusion Failing to let an individual know about the information that others have about them and participate in its handling or use A company using customer call history, without the customer’s knowledge, to shift their order in a queue (i.e. "Your call will be answer in the order [NOT] received") to have the authority of controlling information about yourself who can or can not see. to be from from any interference, and to be able to interact with anyone I want.
Breach of Confidentiality Breaking a promise to keep a person’s information confidential A doctor revealing patient information to friends on a social media website. Having confidential and private information secured and stored away safely from malicious users.
Disclosure Revealing truthful personal information about a person that impacts the ways others judge their character or their security A government agency revealing an individual’s address to a stalker, resulting in the individual’s murder. Data must be kept safe, and users need that information to be seen only by those they authorize.
Exposure Revealing an individual’s nudity, grief, or bodily functions A store forcing a customer to remove clothing revealing a colostomy bag. Freedom of your own information.
Increased Accessibility Amplifying the accessibility of personal information A court making proceeding searchable on the Internet without redacting personal information. A state where one can be sure no one else knows what they are doing
Blackmail Threatening to disclose personal information A dating service for adulters charging customers to delete their accounts. -
Appropriation Using an individual’s identity to serve the aims and interests of another A social media site using customer’s images in advertising Being able to be secure in your information so that none of it gets accessed or leaked by outside sources
Distortion Disseminating false or misleading information about an individual A creditor reporting a paid bill as unpaid to a credit bureau. Privacy refers to an individual’s right to control [..]on. This includes protecting sensitive data from [..], and providing individuals with the ability to access, correct, or delete their PI.
Intrusion Disturbing an individual’s tranquility or solitude An augmented reality game directing players onto private residential property. The right to be let alone,or freedom from interference or intrusion.
Decisional Inference Intruding into an individual’s decision regarding their private affairs A payment processor declining transactions for contraceptives The right to be let alone,or freedom from interference or intrusion.